OpenStudy (anonymous):

**GIVING MEDAL** Write the equation in slope-intercept form. What are the slope and y-intercept? -5x - 12y = 11

OpenStudy (anonymous):

y=(5/12)x+(11/12) m=(5/12), c=(11/12)